OpenStudy (anonymous):

Solve for x:1/3 (2x − 8) = 4 x = 2 x = 6 x = 10 x = 10/3

OpenStudy (callisto):

\[\frac{1}{3}(2x-8)=4\]If this is the question, multiply both sides by 3 first.

OpenStudy (callisto):

What would you get?

OpenStudy (anonymous):

dont get it

OpenStudy (callisto):

Is \(\frac{1}{3}(2x-8)=4\) the equation you need to solve?

OpenStudy (anonymous):

ya thats it

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }\left( 2x-8 \right)=4\] Multiply both sides by 3. It will get rid of the 1/3 and make the equation simple. so: \[\frac{ 1 }{ 3 }(3)(2x-8)=4(3)\] Note that the terms in the parenthesis are being left alone, since they are inside the parenthesis, lol so now we get: 1(2x-8)=12 -> 2x-8=12 PLus 8 on both sides +8 +8 2x=20 divide by 2 on both sides to get the variable alone /2 /2 final answr x=10